# Process-based Indicators of Vulnerability Re-Introducing Code Changes: An Exploratory Case Study

### v1

@@ 3.2.1. Dataset Collection
As current vulnerability datasets do not provide vulnerability reintroducing and fixing commit pairs, we developed a ground truth dataset for ImageMagick that extends existing vulnerability datasets with vulnerability reintroducing commits. We first extracted existing ImageMagick vulnerability commits from BigVul (Fan et al., 2020) and DiverseVul (Chen et al., 2023), two existing and well-studied vulnerability datasets. For our initial case study, we filtered for commits that modified only one file. To filter for commits that specifically reintroduced vulnerabilities, we analyzed subsequent commits to determine whether regressions occurred that modified the initial fix with the SZZ algorithm (Śliwerski et al., 2005). As SZZ solely operates on source code, we also manually evaluated the commit message attached to SZZ identified commits. To further refine our dataset we leveraged Anthropic Claude 3.5 Sonnet (https://claude.ai/) to confirm if each vulnerability reintroducing and fixing commit was accurate. Our prompt to do so is presented in Listing 3.
@@
Our dataset collection resulted in 81 confirmed vulnerability reintroducing-fixing commit pairs for ImageMagick from an initial set of 175 and 225 commits from BigVul and DiverseVul respectively.Of the 81 pairs, 25 originated from BigVul (Fan et al., 2020) and 56 originated from DiverseVul (Chen et al., 2023), however only 76 have associated CVEs. Our dataset is publicly available [^3] https://github.com/anonSubmissionGithub/ProcesseMetric-VulReintroduction.
@@ 3.2.2. CVE Selection Criteria
From the 76 candidate pairs of vulnerability reintroducing commits, we randomly selected four pairs. Our selection criteria were that each pair was identified to resolve a known CVE with at least a score of 8.8 while reintroducing a new vulnerability, that each vulnerability pair is non-overlapping, and each vulnerability had to be identified in a unique year. Our candidate vulnerabilities are presented in Table˜1.
@@ 3.2.3. Software Process Metrics Analysis
For longitudinal computation of process metrics, we utilized the PRIME tool (Synovic et al., 2022). PRIME can compute bus factor, issue density, and issue spoilage among others for any GitHub hosted source code repository with a GitHub issue tracker. For bus factor, we computed the bus factor for the entire project every six months. For issue density and spoilage, we computed the metric per six months, but for each CVE in our case study we computed the metric per week for 20 weeks prior to, during, and 20 after the reintroduced vulnerability persisted in the project.
@@ 3.3. Data Analysis
In this section we discuss an overview of the dataset of 76 candidate vulnerability reintroducing-fixing commit pairs identified in Section˜3.2.1, and our longitudinal analysis of bus factor, issue density, and issue spoilage as it relates to vulnerability reintroduction activities.
@@ 3.3.1. ImageMagick Vulnerabilities per Year:
Table˜2 breaks down the 76 out of 81 identified vulnerability reintroducing commits with CVEs and their average CVSS scores per year. Since 2017, the number of vulnerabilities that are reintroduced into ImageMagick has decreased. Furthermore, the average CVSS score of vulnerabilities that reintroduce vulnerabilities has also decreased. This trend suggests that ImageMagick’s contributors are becoming increasingly effective at identifying, isolating, and mitigating security flaws before they escalate in severity. It also indicates a maturation of the project’s secure development practices, where vulnerability resolution attempts are less likely to result in the reintroduction of high-impact security issues.
@@ 3.4. Bus Factor
ImageMagick’s bus factor per six months metric is presented in Figure˜4. ImageMagick’s has had at least four contributors per six months with as many seven contributing to the project. This guarantees that at any one time, there have been a sufficient number of contributors who can resolve vulnerabilities. It also implies that there is a probability that any any given contributor could reintroduce a vulnerability while attempting to resolve a CVE.
@@ 3.5. Issue Density
Figure 5. ImageMagick’s issue density peaked in 2022 with 28% issue density, but has since dropped to 16.4% as of writing. CVE-2016-4564 saw increasing issue density because the rate of new issues being opened exceeded the maintainers’ closure rate. CVE-2017-16546 and CVE-2018-11625 both experienced density dips due to code deletion and the closing of numerous issues, respectively. CVE-2019-13299 showed increasing density driven by a rise in open issues.
@@
ImageMagick’s and selected CVEs’ issue density is presented in Figure˜5. ImageMagick’s issue density per six months is low, at only 28% of open issues to total project size at its peak in 2020.
@@
CVE-2016-4564: Prior to the vulnerability being reintroduced, issue density was increasing. After the vulnerability was reintroduced, issue density slightly decreased. After the reintroduced vulnerability was fixed, issue density decreased weekly, until rising once more. When evaluating the rate at which the project size was changed — measured in thousands of lines of code per week (i.e., KLOC per week) — during this time frame, we see that it experiences little change. Thus, the issue density of this CVE is reflective of the contributors initially closing issues raised by the community, before the community started contributing more issues than the author’s could close.
@@
CVE-2017-16546: Prior to the vulnerability being reintroduced, issue density was decreasing. After the vulnerability was reintroduced, issue density experienced a wave of increasing issue density, before issue density decreased. After the reintroduced vulnerability was fixed, issue density increased. When evaluating the KLOC per week during this time frame, we see that it dips early in the time frame before returning to normal. Thus, the issue density of this CVE is reflective of code being removed from the project before being added back, resulting in a noticeable spike in issue density at this time.
@@
CVE-2018-11625: Prior to the vulnerability being reintroduced, issue density was increasing. After the vulnerability was reintroduced, issue density continued to rise for a five weeks before sharply decreasing. This followed by another increase in issue density, followed by a decrease, then continuous increasing throughout the remainder of the time frame and after the fix. When evaluating the KLOC per week during this time frame, we see that it increases during this time frame. Thus, the issue density of this CVE is reflective of the project maintainers both modifiying the code and closing issues concurrently during this time frame.
@@
CVE-2019-13299: Prior to and during the vulnerability reintroduction, and after the fix was released, issue density was increasing. When evaluating the KLOC per week during this time frame, we see that it is stable. Thus, the issue density of this CVE is reflective of an increasing amount of issues being opened during this time frame.
Issue density exhibited three distinct trajectories across our four selected CVEs:
  • Declining stability (1/4 CVEs – CVE-2016-4564): Density decreased after the risky fix, implying improved triage and faster issue closure.
  • Fluctuating recovery (2/4 CVEs – CVE-2017-16546, CVE-2018-11625): Density oscillated due to simultaneous code refactoring and issue handling.
  • Persistent growth (1/4 CVEs – CVE-2019-13299): Density continuously increased, indicating sustained issue accumulation and limited closure capacity.
These findings suggest that vulnerability reintroduction is often accompanied by temporary disruptions in issue management efficiency, reflecting the project’s fluctuating capacity to balance bug triage and ongoing development. This observation is particularly interesting and warrants further investigation across larger and more diverse projects.
@@ 3.6. Issue Spoilage
Figure 6. ImageMagick’s issue spoilage peaked in 2023 with 74,657 days of spoilage, but has since dropped to 25 days as of writing. CVE-2016-4564 saw an increase in issue spoilage after the vulnerability reintroduction, dipping only after the corrective fix was published. CVE-2017-16546 and CVE-2018-116 both showed a dip in issue spoilage immediately after vulnerability reintroduction. CVE-2019-13299 experienced increasing issue spoilage, implying that ImageMagick was unable to keep up with closing related issues in a timely manner.
ImageMagick’s and selected CVEs’ issue spoilage is presented in Figure˜6. ImageMagick’s issue spoilage per six months is high but showing improvement, with over 74,657 days of spoiled issues at its peak in 2023, and currently at 25 days of spoiled issues as of writing.
CVE-2016-4564: Prior to and during the reintroduced vulnerabilities time frame, issue spoilage was increasing. After the reintroduced vulnerability was fixed, issue spoilage decreased before before increasing again. This reflects that towards the tail end of the reintroduction period, the project maintainers put in the effort to resolve long open issues before resuming other activities.
CVE-2017-16546: Prior to the vulnerability being reintroduced, issue spoilage was decreasing. After the vulnerability was reintroduced, issue spoilage experienced a wave of increasing issue spoilage, a dip, and then increasing issue spoilage. After the reintroduced vulnerability was fixed, issue spoilage continued to increase. This reflects that during this reintroduced vulnerabilities life, the maintainers were actively resolving issues while conducting other activities.
CVE-2018-11625: Prior to the vulnerability being reintroduced and initially after the introduction, issue spoilage was increasing. During the reintroduction time frame, issue spoilage experienced a wave of increasing issue spoilage, a dip, and then continuous increase throughout the time frame and after. This shows that the maintainers initially closed a significant number of issues, and then continued to keep up with resolving open issues throughout the vulnerabilities reintroduction period.
CVE-2019-13299: Prior to and during the vulnerability reintroduction, and after the fix issue spoilage was increasing. This is reflective of the maintainers not being able to close open issues faster than they were being opened.
Issue spoilage exhibited three distinct patterns across our four selected CVEs:
  • Reactive escalation (1/4 CVEs): Spoilage increased after risky fix
  • Aggressive velocity (2/4 CVEs): Spoilage declined after risky fix
  • Sustained degradation (1/4 CVEs): Spoilage continuously increased
These patterns collectively indicate that fluctuations in issue spoilage correspond to varying levels of responsiveness and process stability during vulnerability reintroduction periods.
@@ 4. Threats To Validity
We report our findings based on an initial study of a subset of vulnerability reintroduction commit pairs from ImageMagick — a well known and embeddedable image processing project. The relatively small sample size of analyzed commti pairs limits the generalizability of our conclusions. Since ImageMagick is primarily written in C/C++, the results may not generalize to projects in other languages or domains. Furthermore, not all open-source projects publicly disclose vulnerabilities to centralized vulnerability databases, potentially limiting the applicability of our approach to the wider open-source community. To mitigate these threats we relied on existing vulnerability datasets (Fan et al., 2020; Chen et al., 2023) to identify projects with identified CVE and CWEs to focus on projects of interest to the software engineering community.
Our methodology to identify vulnerability reintroducing and fixing commit pairs leveraged Anthropic Claude 3.5 Sonnet LLM to confirm manually identified commit pairs from the BigVul vulnerability dataset (Fan et al., 2020). We also leveraged the LLM to autonomously identify commit pairs from the DiverseVul vulnerability dataset (Chen et al., 2023). This decision was driven by the significant time and complexity involved in conducting a full manual analysis.
Furthermore, our leveraged bus factor metric (Cosentino et al., 2015) measures the aggregate number of unique contributors to a project, and does not take into account who contributes to individual directories or files. Thus, while our findings show that at least four contributors are working on ImageMagick at any one time, it does not enable the evaluation of the engineering activity of each contributor. To mitigate this, we do not focus on the engineering contributions of any one contributor, but rather looking at the wholistic effort of the contributing team.
@@ 5. Discussion and Future Implications
This initial case study on ImageMagick demonstrates the value of analyzing SEP metrics to understand vulnerability reintroduction patterns. Our findings reveal potential insight on how bus factor, issue density, and issue spoilage correlate with vulnerability reintroduction activities. These findings highlight several directions for future research.
Expanding the Scope Beyond ImageMagick First, we will expand the scope of analysis to multiple open-source systems across diverse domains and programming languages. By extending our methodology to diverse projects, we want to assess whether there is any general correlation between SEP and vulnerability reintroduction. This large-scale analysis will allow us to determine if certain socio-technical configurations are consistent among projects.
Developing Predictive Models for Risky Fixes The overarching objective of this research is to enable actionable prediction of vulnerability reintroduction. By integrating longitudinal process metrics with code-level features and commit metadata, we aim to develop machine learning classifiers capable of identifying vulnerability-fixing commits that exhibit a high likelihood of reintroducing new security issues at the time they are proposed. Such predictive systems could inform automated regression testing and security-focused quality assurance workflows.
Investigating Pull Request Dynamics Future work will examine how code review processes influence vulnerability reintroduction. We will analyze metrics such as pull request spoilage, review latency, reviewer-to-contributor ratios, and comment volume to assess their correlation with reintroduction risk. Understanding these dynamics will help determine whether certain review practices, such as rushed approvals, limited reviewer diversity, or insufficient review iterations, are associated with higher vulnerability reintroductions.
@@ 6. Conclusion
This work aims to highlight the critical role of software engineering process metrics in understanding and mitigating the risk of vulnerability reintroduction. By shifting the focus from file-level vulnerability prediction to commit-level analysis of security fixes, we underscore the importance of detecting risky changes not only at the moment they are introduced, but also across long sequence of changes that capture how vulnerabilities evolve and re-emerge over time. Our approach emphasizes that vulnerability reintroduction is rarely the result of a single isolated action, but instead emerges from cumulative sequences of development activities and socio-technical conditions.

### v2

@@ 3.2.1. Dataset Collection
Existing vulnerability datasets do not track the full commit history of vulnerability reintroduction and subsequent fixes. To address this, we constructed a ground truth dataset by extending existing datasets that capture ImageMagick vulnerability records. We first extracted existing ImageMagick vulnerability commits from two exemplar datasets: BigVul (Fan et al., 2020) and DiverseVul (Chen et al., 2023). For our initial case study, we filtered for commits that modified only one file.
To filter for commits that specifically reintroduced vulnerabilities, we analyzed subsequent commits from the identified set of ImageMagick vulnerability fixing commits. To determine whether regressions occurred that reintroduced vulnerabilities, we leveraged the SZZ algorithm, a widely known approach for identifying defect inducing source code (Śliwerski et al., 2005). For each vulnerability fixing commit, we used SZZ to trace the modified lines of code through the version control history to identify all modifying commits, thereby identifying the proper corrective fix if available. Additionally, we also manually evaluated the commit message attached to each SZZ identified corrective commits to determine if the commit resolved a reintroduced vulnerability.
To further refine our dataset and ensure the correctness of each vulnerability reintroducing commit, we leveraged five LLMs: Claude-3.5-Sonnet (Anthropic), GPT-4o (OpenAI), DeepSeek Chat (DeepSeek AI), and Grok-2 (xAI). Based on our sample-level evaluation, we selected Claude as the primary model for constructing the vulnerability reintroducing dataset. All Claude-generated labels were then manually reviewed as a second validation step to ensure accuracy. Our prompt to do so is presented in Listing 3.
@@
Our dataset collection resulted in 81 confirmed vulnerability reintroducing-fixing commit pairs for ImageMagick from an initial set of 175 and 225 commits from BigVul and DiverseVul respectively. Of the 81 pairs, 25 originated from BigVul (Fan et al., 2020) and 56 originated from DiverseVul (Chen et al., 2023), however only 76 have associated CVEs. Our dataset is publicly available (Shimmi et al., 2026)
@@ 3.2.2. CVE Selection Criteria
From the 76 candidate pairs of vulnerability reintroducing commits identified in Section˜3.2.1, we randomly selected four pairs. Our selection criteria were that each pair was identified to resolve a known CVE with at least a score of 8.8 while reintroducing a new vulnerability, that each vulnerability pair is non-overlapping, and each vulnerability had to be identified in a unique year. Our candidate vulnerabilities are presented in Table˜1.
@@ 3.2.3. Software Process Metrics Analysis
For longitudinal computation of process metrics, we utilized the PRIME tool (Synovic et al., 2022). PRIME can compute bus factor, issue density, and issue spoilage among others. For bus factor, we computed the bus factor for the entire project every six months. For issue density and spoilage, we computed the metric per six months, but for each CVE in our case study, we computed the metric per week for 20 weeks prior to, during, and 20 after the reintroduced vulnerability persisted in the project.
@@ 3.3. Data Analysis
In this section, we discuss an overview of the dataset of 76 candidate vulnerability reintroducing-fixing commit pairs identified in Section˜3.2.1, and our longitudinal analysis of bus factor, issue density, and issue spoilage as it relates to vulnerability reintroduction activities.
@@ 3.3.1. ImageMagick Vulnerabilities per Year:
Table˜2 breaks down the 76 identified vulnerability reintroducing commits with CVEs and their average CVSS scores per year. Since 2017, the number of vulnerabilities that are reintroduced into ImageMagick has decreased. Furthermore, the average CVSS score of vulnerabilities that reintroduce vulnerabilities has also decreased. This suggests that ImageMagick’s contributors are becoming increasingly effective at identifying, isolating, and mitigating security flaws prior to including them in the project. It also indicates a maturation of the project’s secure development practices, where vulnerability resolution attempts are less likely to result in the reintroduction of high-impact security issues.
@@ 3.4. Bus Factor
ImageMagick’s bus factor per six months metric is presented in Figure˜4. ImageMagick has at least four—and sometimes as many as seven—contributors during a six-month period. This guarantees that at any period, there have been a sufficient number of contributors who can introduce and resolve vulnerabilities.
@@ 3.5. Issue Density
Figure 5. Since peaking at 28% in 2022, issue density has declined to 16.4%. Fluctuations are tied to specific vulnerabilities: CVE-2016-4564 and CVE-2019-13299 saw density increases driven by rising open-issue rates. Conversely, CVE-2017-16546 and CVE-2018-11625 experienced declines due to code deletion and high issue-closure rates, respectively.
@@
ImageMagick’s and selected CVEs’ issue density is presented in Figure˜5. ImageMagick’s issue density per six months is low, at only 28% of open issues to total project size at its peak in 2020. Our findings suggest that vulnerability reintroduction is often accompanied by temporary disruptions in issue management efficiency, reflecting the project’s fluctuating capacity to balance bug triage and ongoing development.
@@
CVE-2016-4564 Prior to the fix for CVE-2016-4564 in which a vulnerability was reintroduced, issue density was increasing. After the vulnerability was reintroduced, issue density decreased until after the corrective fix was applied. When evaluating the rate at which the project size was changed — measured in thousands of lines of code per week (i.e., KLOC per week) — during this time frame, we see that it experiences little change. Thus, the issue density of this CVE is reflective of contributors initially closing issues raised by the community, before the community started contributing more issues than the author’s could close.
@@
CVE-2017-16546 Prior to the fix for CVE-2017-16546 in which a vulnerability was reintroduced, issue density was decreasing. After the vulnerability was reintroduced, ImageMagick underwent a wave of increasing and decreasing issue density. However, after the proper corrective fix was applied, issue density increased. When evaluating the KLOC per week during this time frame, we see that it initially dips before increasing to it’s initial size. Thus the issue density of this CVE is reflective of both a shrinking code size and closure of open issues.
@@
CVE-2018-11625 Prior to the fix for CVE-2018-11625 in which a vulnerability was reintroduced, and after the proper corrective fix was applied, ImageMagick experienced increasing issue density. During the existence of the reintroduced vulnerability, issue density underwent a wave of increasing and decreasing issue density, until issue density continuously rose post the proper corrective fix being applied. When evaluating the KLOC per week during this time frame, we see that it increases during this time frame, reflecting that the project maintainers were modifying the code and closing issues concurrently during this time frame.
@@
CVE-2019-13299 Prior to the fix for CVE-2019-13299 in which a vulnerability was reintroduced, and after the proper corrective fix was applied, issue density was increasing. When evaluating the KLOC per week during this time frame, we see that it remains stable. Thus, the issue density of this CVE is reflective of an increasing amount of issues being opened during this time frame.
@@ 3.6. Issue Spoilage
Figure 6. After a 2023 peak of 74,657 days, spoilage plummeted to just 25 days. Trends varied by vulnerability: CVE-2016-4564 and CVE-2019-13299 saw rising spoilage following reintroduction—indicating a struggle to resolve issues timely—until corrective patches were applied. In contrast, CVE-2017-16546 and CVE-2018-11625 experienced immediate spoilage dips following vulnerability reintroduction.
ImageMagick’s and selected CVEs’ issue spoilage is presented in Figure˜6. ImageMagick’s issue spoilage per six months is high but showing improvement, with over 74,657 days of spoiled issues at its peak in 2023, and currently at 25 days of spoiled issues as of writing. Our findings indicate that fluctuations in issue spoilage correspond to varying levels of responsiveness and process stability during vulnerability reintroduction periods.
CVE-2016-4564 Prior to the fix for CVE-2016-4564 in which a vulnerability and thereafter, issue spoilage was increasing. After the proper corrective fix was applied, issue spoilage decreased before increasing again. This reflects that towards the tail end of the reintroduced vulnerability’s existence, project maintainers resolved long open issues before resuming other activities.
CVE-2017-16546 Prior to the fix for CVE-2017-16546 in which a vulnerability was reintroduced, issue spoilage was decreasing. After the vulnerability was reintroduced, ImageMagick experienced increasing issue spoilage, a dip, and then increasing issue spoilage after the proper corrective fix was applied. This reflects that during this reintroduced vulnerabilities life, the maintainers were actively resolving issues while conducting other activities.
CVE-2018-11625 Prior to the fix for CVE-2018-11625 in which a vulnerability was reintroduced and upon reintroduction, issue spoilage was increasing. While the reintroduced vulnerability was active, ImageMagick experienced increasing issue spoilage, a dip, and then continuous increase after proper resolution. This reflects maintainers initially closing a significant number of issues, and then continuing to resolve issues throughout the reintroduced vulnerabilities existence.
CVE-2019-13299 Prior to the fix for CVE-2019-13299 in which a vulnerability was reintroduced, and after the proper corrective fix was applied, issue spoilage was increasing. This is reflective of the maintainers not being able to close open issues faster than they were being opened.
@@ 4. Threats To Validity
We report our findings based on an initial study of a subset of vulnerability reintroduction commit pairs from ImageMagick — a well known image processing project. The relatively small sample size of analyzed commit pairs limits the generalizability of our conclusions. Since ImageMagick is primarily written in C/C++, the results may not generalize to projects in other languages or domains. Furthermore, not all open-source projects publicly disclose vulnerabilities to centralized vulnerability databases, potentially limiting the applicability of our approach to the wider open-source community. To mitigate these threats we relied on existing vulnerability datasets (Fan et al., 2020; Chen et al., 2023) to identify projects with identified CVE and CWEs to focus on projects of interest to the software engineering community.
Our methodology to identify pairs of vulnerability reintroducing and fixing commits leverages the Anthropic Claude 3.5 Sonnet LLM to validate manually identified pairs from the BigVul vulnerability dataset (Fan et al., 2020). Additionally, it autonomously identifies commit pairs from the DiverseVul vulnerability dataset (Chen et al., 2023). This decision was driven by the significant time and complexity involved in conducting a full manual analysis.
Furthermore, our leveraged bus factor metric (Cosentino et al., 2015) measures the aggregate number of unique contributors to a project, and does not take into account who contributes to individual directories or files. Thus, while our findings show that at least four contributors are working on ImageMagick at any one time, it does not enable the evaluation of the engineering activity of each contributor. To mitigate this, we focus the holistic effort of the team, rather than specific engineers.
@@ 5. Discussion and Future Implications
This initial case study on ImageMagick demonstrates the value of analyzing software engineering process metrics to understand vulnerability reintroduction patterns. Our findings reveal potential insight on how bus factor, issue density, and issue spoilage correlate with vulnerability reintroduction activities. These findings highlight several directions for future research.
Expanding Beyond ImageMagick We intend to expand the scope of analysis to multiple open-source systems across diverse domains and programming languages. By extending our methodology to diverse projects, we want to assess whether there is any general correlation between software engineering processes and vulnerability reintroduction. This large-scale analysis will allow us to determine if certain socio-technical patterns are consistent among projects.
Developing Predictive Models for Risky Fixes The overarching objective of this research is to enable actionable prediction of vulnerability reintroduction. Future work will develop machine learning classifiers trained on longitudinal process metrics, code-level features, and commit metadata to predict vulnerability-fixing commits that are likely to reintroduce new security issues before they are proposed. Such predictive systems could inform automated regression testing and security-focused quality assurance workflows.
Investigating Pull Request Dynamics Future research will explore how code review processes impact vulnerability reintroduction. By analyzing metrics pertaining to pull requests (e.g., pull request spoilage, review latency, reviewer-to-contributor ratios, comment volume) we aim to evaluate their relationship with reintroduction risk. This analysis will help identify specific pull request practices that contribute to higher vulnerability reintroductions.
@@ 6. Conclusion
To our knowledge, this is the first work to highlight the critical role of software engineering process metrics in understanding and mitigating the risk of vulnerability reintroduction. By shifting focus from file-level vulnerability prediction to analyzing sequences of security fixing commits, we underscore the importance of detecting vulnerability enabling engineering activities longitudinally. Our approach emphasizes that vulnerability reintroduction is rarely the result of a single isolated action, but instead emerges from cumulative sequences of engineering activities and socio-technical conditions.
Acknowledgements.
The work in this paper was partially funded by the Office of Naval Research (ONR) (Grant#: G2A62826).